Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability - 100 Numbered Balls

Expert replies
by evansbd » Wed Nov 26, 2008 8:56 am
A box contains 100 balls, numbered from 1 to 100. If three balls are selected at random and with replacement from the box, what is the probability that the sum of the three numbers on the balls selected from the box will be odd?

A) 1/4 B) 3/8 C) 1/2 D) 5/8 E) 3/4


I wanted to see how others approached this problem. Luckily I've posted and practiced many many probability questions in this forum, so I got this one right. However I did it by inspection, so if others could answer it I can see if my approach is good in general or simply worked for this particular question.
Join the discussion
Source: — Problem Solving |

by mals24 » Wed Nov 26, 2008 9:07 am
Ok number of ways the sum of 3 digits can be odd:

E+E+O
O+O+O

(E - Even, O - Odd)

E E O can further be arranged in 3 ways.

So total probability = (1/2*1/2*1/2)*3 + (1/2*1/2*1/2) = 3/8+1/8 = 4/8 = 1/2.

IMO C.
Join the discussion

by gmataug08 » Wed Nov 26, 2008 9:11 am
there are 50 odd balls and 50 even balls

and the possible results are either odd or even

possible combination for ODD result :

O + E + E & O + O + O

possible combination for even result :

E + O + O & E + E + E

hence the probability would be 1/2

imo, answer C .

what is OA?
Join the discussion

by dmateer25 » Wed Nov 26, 2008 9:11 am
Here is the approach I took:

Here are the possible outcomes for the balls:

E+E+E = Even
E+E+O = Odd
E+O+O = Even
E+O+E = Odd
O+E+E = Odd
O+E+O = Even
O+O+E = Even
O+O+O = Odd


4 possibilities it will be even and 4 that it will be odd

so 4/8 that it will be odd

1/2 that it will be odd
Join the discussion

by muzali » Wed Nov 26, 2008 9:16 am
Alternatively, _ _ _ (need to fill w/ E or O)

this is like a coin flip question, so total ways = 2^n = 2^3 = 8
Half of these will add to even and the other half to odd, so probability = 1/2
Join the discussion

by cramya » Wed Nov 26, 2008 9:25 am
3 numbers to choose and 2 choices(EVEN OR ODD) for each number

2*2*2 = 8 different outcomes

OOE
OEO
EOO
EEE
OOO
EEO
EOE
EOO

3/8 cases ithe sum will be odd
Join the discussion

by dmateer25 » Wed Nov 26, 2008 9:39 am
cramya wrote:3 numbers to choose and 2 choices(EVEN OR ODD) for each number

2*2*2 = 8 different outcomes

OOE
OEO
EOO
EEE
OOO
EEO
EOE
EOO

3/8 cases ithe sum will be odd
OOE=even
OEO=even
EOO =even
EEE =even
OOO =odd
EEO =odd
EOE =odd
EOO ( should be OEE not EOO so this would be odd)
Join the discussion

Probability - 100 numbered balls

by evansbd » Sat Nov 29, 2008 12:16 pm
The OA is 1/2.

Everyone has the right idea. We know that the probability of even or odd is 1/2.
dmateer25 wrote:Here is the approach I took:

Here are the possible outcomes for the balls:

E+E+E = Even
E+E+O = Odd
E+O+O = Even
E+O+E = Odd
O+E+E = Odd
O+E+O = Even
O+O+E = Even
O+O+O = Odd
The OA is 1/2.

Since the order doesn't matter (with replacement), you can see here that 1/2 of your options will give you even or odd. If the question was without replacement then you would need to do some calculations. I guess either way you get the right answer by doing it like a coin flip but inspection, if you can do it, saves time.

What would be the answer w/o replacement?
Join the discussion

Re: Probability - 100 numbered balls

by abcdefg » Tue Jul 14, 2009 6:29 pm
I'm confused by what you mean when you say "order doesn't matter." If order doesn't matter then E-E-O and O-E-E are the same thing. But then why are we listing them out to be different possible outcomes?
evansbd wrote:The OA is 1/2.

Everyone has the right idea. We know that the probability of even or odd is 1/2.
dmateer25 wrote:Here is the approach I took:

Here are the possible outcomes for the balls:

E+E+E = Even
E+E+O = Odd
E+O+O = Even
E+O+E = Odd
O+E+E = Odd
O+E+O = Even
O+O+E = Even
O+O+O = Odd
The OA is 1/2.

Since the order doesn't matter (with replacement), you can see here that 1/2 of your options will give you even or odd. If the question was without replacement then you would need to do some calculations. I guess either way you get the right answer by doing it like a coin flip but inspection, if you can do it, saves time.

What would be the answer w/o replacement?
Join the discussion

Re: Probability - 100 Numbered Balls

by doclkk » Wed Jul 15, 2009 9:57 am
evansbd wrote:A box contains 100 balls, numbered from 1 to 100. If three balls are selected at random and with replacement from the box, what is the probability that the sum of the three numbers on the balls selected from the box will be odd?

A) 1/4 B) 3/8 C) 1/2 D) 5/8 E) 3/4


I wanted to see how others approached this problem. Luckily I've posted and practiced many many probability questions in this forum, so I got this one right. However I did it by inspection, so if others could answer it I can see if my approach is good in general or simply worked for this particular question.
Should be C

Why can't your thinking just be half the numbers from 1-100 are odd ? Forget about sum for a minute. What's the probably of any ball that's picked will be odd right? Half the balls are odd. Half the balls are even. nuff said? I don't see the need for all of the E+O+E etc exercise.
Join the discussion

by aircraft » Wed May 15, 2013 3:14 pm
abcdefg wrote:I'm confused by what you mean when you say "order doesn't matter." If order doesn't matter then E-E-O and O-E-E are the same thing. But then why are we listing them out to be different possible outcomes?
evansbd wrote:The OA is 1/2.

The OA is 1/2.

Since the order doesn't matter (with replacement), you can see here that 1/2 of your options will give you even or odd. If the question was without replacement then you would need to do some calculations. I guess either way you get the right answer by doing it like a coin flip but inspection, if you can do it, saves time.
I'm actually also struggling with that question: why isn't EOO treated the same way as OOE if order doesn't matter?
Join the discussion